# MiniScope: Automated UI Exploration and Privacy Inconsistency Detection of MiniApps via Two-phase Iterative Hybrid Analysis

### v1

@@ 3. Motivating Example
We provide a running example showing how MiniApp collects and uses privacy in Figure 2, which consists of three key stages: (1) Cross-package page routing. The user triggers cross-package page routing via <navigator> (❶) and <botton> (❹❺) widgets in pages/myInfo (main package), respectively navigating to pages/takePhoto (main package) and pages/checkID (subpackage dynamically loaded); (2) Event-driven callback execution. The user triggers binding events in pages/takePhoto (❷) and pages/checkID (❻) through UI interactions, and the logic layer (JavaScript) executes the event-driven callback functions takePhoto and onInput respectively; (3) Sensitive data flow propagation. In the asynchronous function callbacks, the sensitive data flow further completes the propagation (❸❼). Notably, in this example, due to the lack of sufficient page routing and binding events, the pages/demo in the sub-package and the startRecord function within it are considered to be unreachable, which could potentially leads to FPs in previous DFA-only approaches (Wang et al., 2023b; Meng et al., 2023; Li et al., 2023).
@@
Figure 2. Code snippet of cross-file/cross-package privacy practices. Subfigure (a) and (b) illustrate cross-file privacy practices in the initially loaded main package, subfigure (a) and (c) represent cross-subpackage privacy practices (i.e., FNs in previous static-only approaches), and subfigure (d) demonstrates dead code without entry points (i.e., FPs generated in previous DFA-only approaches).
@@ 3.1. Challenges for Hybrid Analysis of MiniApps
Considering the unique nature of MiniApps, which relies on Web technologies and integrates with the capabilities of native apps, the hybrid analysis of MiniApps primarily faces two key challenges:
Challenge#1: Complexity of cross-file/package callback control flow resolution. In MiniApps, a prominent challenge emerges from the complexity in analyzing control flow, particularly due to the widespread practice of reusing callback functions across different files and subpackages. This aspect of MiniApp development present a unique scenario, which is not directly analogous to existing techniques in Android/iOS environments. Developers often define callback function logic in shared JavaScript files and subsequently import these callbacks into various pages and subpackages of the MiniApp. This results in a complex, intertwined control flow graph of cross-file and cross-(sub)package callbacks, significantly complicating the process of tracing user interactions and understanding event-driven behaviors.
Challenge#2: Inadequate in locating WeChat-specific components in the UI widgets tree The distinctive components within WXML introduce unique challenges in the realm of UI exploration. These challenges primarily stem from the limitations of traditional testing tools. Designed specifically to align with the WeChat ecosystem, WXML components diverge significantly from the typical Document Object Model (DOM) structures or native UI frameworks familiar to conventional tools. This divergence renders tools like UIAutomator ineffective in identifying and locating these components directly from the GUI widgets tree. This limitation necessitates the development or adaptation of specialized testing tools and methodologies that can cater to the unique structure and behavior of WXML components, ensuring effective and efficient testing of MiniApps.
@@ 3.2. Insights for Potential Solutions
Upon evaluating the complexities inherent in MiniApps, particularly considering the aforementioned challenges, we distill two refined insights for potential solutions.
Insight#1: Constructing fine-grained and comprehensive topological structures for MiniApps. To address Challenge#1, the primary focus is on constructing a detailed and comprehensive topological structure of MiniApps. This structure would facilitate a more nuanced analysis across files and subpackages, supporting an in-depth control (❶❷) and data flow (❸) analysis. Notably, integrating callback control flow analysis within this framework is crucial for dynamic UI exploration. It could also potentially be a starting point for static data-flow analysis, thereby reducing false positives in static analysis. This insight underscores the need for a multifaceted approach, blending fine-grained structural mapping with advanced flow analysis, to provide a robust foundation for page transition analysis and automated UI exploration.
Insight#2: Enhancing existing dynamic testing tools with fuzzy matching for platform-specific components In response to Challenge#2, the strategy involves augmenting existing dynamic testing tools to better handle platform-specific MiniApp widgets. This enhancement would be achieved by implementing a fuzzy matching mechanism based on key component attributes like name, type, and text. By establishing an accurate correlation between static WXML component and dynamic UI Widget Tree elements, we can significantly improve the identification of MiniApp-specific widgets during dynamic runtime. This approach not only ensures the correct recognition of unique MiniApp components but also enhances the overall effectiveness of dynamic testing tools. This insight highlights the importance of adapting testing methodologies to the specificities of the MiniApp platform, leveraging both static and dynamic analysis attributes for comprehensive testing coverage.
@@ 4. Methodology
In response to the two challenges, we propose MiniScope, an automated UI exploration and privacy inconsistency analysis framework based on two-phase iterative hybrid program analysis, as shown in Figure 3. The framework consists of three key modules: MiniApp Dependency Graph Generator (§4.1), Directed UI Explorer (§4.2), and Privacy Inconsistency Detector (§4.3), which further includes a policy analyzer and a privacy practice monitor, thereby facilitating flow-to-policy cross-validation. Our methodology unfolds in two distinct phases, each involving a round of static analysis followed by a round of UI exploration. The entire workflow is structured as follows:
Main Package Analysis (Phase One). The initial phase of MiniScope is dedicated to dynamically loading subpackages through UI exploration to obtain the full package. Firstly, MiniScope initializes the MiniApp Dependency Graph (MDG) for the main package. Subsequently, based on the guidance of MDG, the Directed UI Explorer employs a subpackage-directed breadth-first traversal strategy to explore subpackage pages, thus obtaining the full package of the MiniApp.
Full Package Analysis (Phase Two). In the second phase, MiniScope tackles the precise identification of privacy practices based on the analysis of merged full package. Mirroring the approach of the previous phase, we commence with constructing the complete MDG through static analysis. During the UI exploration stage, MiniScope adopts a privacy-practice-directed depth-first strategy, which is instrumental in conducting a runtime exploration of sensitive behaviors, allowing for a nuanced understanding of the MiniApp’s privacy practices.
@@ 4.1.1. Analyzing UI State Transitions
UTG construction. During UTG construction, MiniScope analyzes the rendering and logic layers separately: (1) For the rendering layer, MiniScope focuses on the <navigator> widgets in the WXML file responsible for page navigation, with the open-type attribute signifying the routing method. (2) For the logic layer, MiniScope chiefly evaluates JavaScript code related to page routing APIs. For instance, wx.navigateTo adds the current page to the page stack and navigates to a new page. By examining five methods in the <navigator> widget and five types of page routing APIs in the logic layer as shown in Table 2, MiniScope can delineate all possible page transitions in the parsing function $\lambda_{U}$ and statically construct the UI state transition graph $G_{U}$.
Figure 4. A simplified UTG example of Express MiniApp. Corresponding to Figure 2, in Event 1, the user triggers ontap event of <navigator>, leading to a transition from pages/myInfo to pages/takePhoto; in Event 2, the user triggers ontap event of <button text="manual input">, and the corresponding event handler callback function naveToCheckID is executed, naveToCheckID callback function, resulting in a transition from pages/myInfo to pages/checkID.
A UTG example. Figure 4 illustrates a simplified UTG example of the Express MiniApp. In this example, the set $V_{U}$ contains three pages: “pages/myInfo/index”(A), “pages/takePhoto/index”(B), and texttt“pages/checkID/index”(C). In the MiniApp framework, all page widgets must be registered in the WXML file. We determine page states $V_{U}$ by traversing through these widgets. For example, PageA consists of six <navigator> widgets, and when we click on the <navigator url=“/pages/takePhoto/index”>, it will navigate to the corresponding page, making the widget a transition condition of the set $E_{U}$ from PageA to PageB.
@@ 4.1.2. Analyzing Callback Control Flows
Determining CCFG entry point. Prior to constructing, it is necessary to determine the CCFG entry points which are event-driven callback functions. This is accomplished by performing cross-language analysis on rendering layer widgets and their associated event handler callback functions. User interactions with these widgets trigger the corresponding event handling logic, activating the associated control flows within the MiniApp. As an example shown in Figure 2 (❷), a widget with “bindtap = takePhoto” attribute will bind to the Tap event, and execute the takePhoto function when a Tap event is triggered. Furthermore, the page lifecycle callbacks can also serve as CCFG entry points as they are tied to specific lifecycle events, executing relevant logic when these events occur.
CCFG construction. From the identified entry points, MiniScope constructs the rest of the function call chain within the JavaScript language to obtain a complete CCFG. Considering the highly customizable, framework-based nature of MiniApps, a tailored cross-file/package callback control flow analysis approach is necessary. Function definitions within MiniApp pages resemble class method definitions, as they’re all defined within the Page() object instance. During initialization, developers provide all lifecycle callback functions, event handler functions, and custom functions to the Page object as attribute methods. Method intercommunication requires the “this” keyword, referring to the current Page object. Notably, some commonly used callback functions can be defined in shared JavaScript files and then imported across various pages. Leveraging these observations, we design a cross-file callback control flow analysis algorithm based on the JavaScript Abstract Syntax Tree (AST). This algorithm takes into account the unique patterns of function usage and invocation in MiniApps, including the shared use of callback functions across different files and subpackages.
@@
Construction algorithm. The main steps of cross-file CCFG construction are described in Algorithm 1. The algorithm takes inputs including the AST $G_{A}$ of the current page, a set of $pageMethods$, a set of $entryPoints$, and a set of $sensiAPIs$. The output of the algorithm is the complete CCFG $G_{C}$. In particular, the algorithm consists of the following three key components:
$\bullet$ Function getImportedMethods(): Iterates through all nodes in the AST. If a node is an $ImportDeclaration$, the $specifier$ of the node is appended to the $pageMethods$ list (line 3-5).
$\bullet$ Function getCCFG(): Iterates through the children of $func$. If a child node is a $CallExpression$, it retrieves the callee of the node (line 11-12). If the callee is a $MemberExpression$ and its left child is a $ThisExpression$ and its right child is in $pageMethods$, it adds an edge to $G_{C}$ and recursively calls $getCCFG$ for this node (line 13-18). If the callee is in $sensiAPIs$, it also adds an edge to $G_{C}$ (line 20-22).
$\bullet$ Main Process: Initializes $pageMethods$ by calling $getImportedMethods$ with the AST $G_{A}$ and an initial list of $pageMethods$ (line 26). Then iterates through the children of the AST $G_{A}$. If a child is a $CallExpression$, it gets the callee of the node (line 28-29). If this node is in $entryPoints$, it calls $getCCFG$ with the node, $pageMethods$, and an initial CCFG $G_{C}$ (line 30-32). Finally, it returns the constructed CCFG $G_{C}$.
Input: AST $G_{A}$ of the page, a set of $pageMethods$, a set of $entryPoints$, a set of $sensiAPIs$
Output: CCFG $G_{C}$ of the page
1 Function getImportedMethods($G_{A},pageMethods$):
@@
4                   $pageMethods.append(node.specifier)$
@@
6
@@
10Function getCCFG($func$, $pageMethods$, $G_{C}$):
11       for $node\in func.children$ do
@@
13                   $callee\leftarrow node.callee$
14                   if $callee$ is $MemberExpression$ then
15                         if $callee.leftchild$ is $ThisExpression$ and $callee.rightchild\in pageMethods$ then
16                               $G_{C}.addEdge(func,node)$
17                               getCCFG($node$, $pageMethods$, $G_{C}$)
18
@@
20
@@
24            if $callee\in sensiAPIs$ then
25                   $G_{C}.addEdge(func,node)$
26
@@
[…]
33 for $node\in G_{A}.children$ do
34       if $node$ is $CallExpression$ then
35             $callee\leftarrow node.callee$
36             if $node\in entryPoints$ then
37                   getCCFG($node$, $pageMethods$, $G_{C}$)
38             end if
39
40       end if
41
42 end for
return $G_{C}$
Algorithm 1 Cross-file CCFG Construction

### v3

@@ 3. Motivating Example
To illustrate the challenges of identifying the privacy practice of MiniApps, we provide a running example in Figure 2, showing how MiniApp collects and uses privacy. In the given MiniApp, (a)(b)(c) are part of the main package, which is initialized during a cold start, while (d)(e)(f) reside within the subpackage that is loaded on-demand when accessed. Since the subpackage code is only available during dynamic access, the static-only analysis may overlook privacy behaviors in (d). Note that the privacy practices exhibited in (b)(c)(d) are potentially triggered by users, whereas the privacy behaviors in (e)(f) become dead code due to unused functions and orphaned pages, and thus will not be invoked. This implies that without UI state transition and callback control flow analysis, the privacy behaviors in (e)(f) might be incorrectly flagged as violations of the privacy policy, which is clearly a false positive. In the following, we describe this process in seven steps.
@@
Figure 2. Code snippet of cross-file/cross-package privacy practices. Subfigure (a), (b), and (c) illustrate cross-file privacy practices in the initially loaded main package, subfigure (a) and (d) represent cross-subpackage privacy practices (i.e., FNs in previous static-only approaches, but identified by MiniScope), and subfigure (e) and (f) demonstrates dead code (unused function and orphaned page, respectively) without entry points (i.e., FPs generated in previous DFA-only approaches, but excluded by MiniScope).
[…]
Main Package Internal Page Routing (WXML $\Rightarrow$ WXML). Page routing facilitates navigation between different views, enabling analysis to determine user-interactable pages. In pages/myInfo, the <navigator> component directs the user to data:takePhotoPath when clicked, which is set to ‘‘pages/takePhoto/takePhoto’’ in the JavaScript logic.
Dynamic-defined Event Handler Callback Control Flow (WXML $\Rightarrow$ JS). Event handler callback control flow is essential for identifying the program logic triggered by user actions. In pages/takePhoto, the <button> widget is bound to the event handler callback function onShutterTap, which is triggered when the button is tapped. Unlike statically defining the callback within the Page object, the callback function onShutterTap is dynamically defined by util.init in onReady.
Cross-page Function Reuse Control Flow (JS $\Rightarrow$ JS). Cross-page function reuse is a critical aspect of maintaining code modularity and efficiency within a MiniApp. In pages/takePhoto, the event handler function onShutterTap is dynamically defined by the function init from pages/util/util.js, which is called with the current page context using this. This is achieved through context binding, ensuring that the onShutterTap has access to the correct page context when invoked.
Data Transmission through Context Binding (JS $\Rightarrow$ JS). In pages/takePhoto, the function takePhoto invokes util.onShutterTap and passes the current page context (this) as an argument to the parameter pageContext. This binding allows the function util.onShutterTap to interact directly with the page’s data and state. Within the function onShutterTap, wx.createCameraContext is invoked to initiate the photo-taking process. Upon successfully capturing an image, the callback success is triggered, and the resulting temporary file path (res.tempImagePath) is passed back to the page via pageContext.setData, setting the property imagePath to the file path.
Main-to-Subpackage Page Routing (JS $\Rightarrow$ WXML). The navigation from pages/myInfo to pages/checkID is triggered by navToCheckID, which utilizes wx.navigateTo to specify the target path. Upon issuing the navigation command, the WeChat dynamically downloads the subpackage code and stores it in the local file system (com.tencent.mm/MicroMsg/hash/appbrand/pkg/ in Android). This dynamic loading ensures that the subpackage code is only fetched and loaded into memory when the user navigates to the corresponding page, thereby reducing the initial load time and memory usage of the MiniApp.
Data Transmission through Event Propagation (WXML $\Rightarrow$ JS). Data transmission through event propagation is a fundamental concept in MiniApp development, facilitating the flow of user input from WXML to the underlying JavaScript logic. Here, the bindinput attribute of the <input> widget binds the event:input to the handler callback function onInput. The function onInput receives the event object as an argument, which contains the detailed user input value event.detail.value. This value is then used to construct a POST request to the server, with the parameter idCard set to the user input.
@@ 3.1. Challenges for Hybrid Analysis of MiniApps
Considering the unique nature of MiniApps, which relies on Web technologies and integrates with the capabilities of native apps, the hybrid analysis of MiniApps primarily faces three key challenges:
Challenge#1: Dynamic Data Binding across Different Languages in UI State Transition Analysis. In MiniApps, dynamic data binding and page routing create complex interactions between JavaScript and WXML. This complexity arises from the need to accurately track how data updates (target page path) propagate through the WXML components and JavaScript logic codes. Such an example is the step ❶ and ❷ in Figure 2. Unlike Android where transitions are often explicitly defined through Intent objects and are relatively straightforward to trace, MiniApp involves a more intricate process due to the dynamic nature of data binding and cross-language interactions. Tracking and analyzing this type of UI state transition is challenging.
Challenge#2: Complexity of Cross-File Callback Control Flow Resolution. Another prominent challenge emerges from the complexity of analyzing control flow, particularly due to the context binding when reusing callback functions across different files and subpackages. In MiniApps, callback functions often rely on the context (this) in which they are executed. Such an example is the step ❸, ❹ and ❺ in Figure 2. Unlike traditional Android or iOS environments, where the scope and lifecycle of callbacks are more contained and predictable, MiniApps allow for the dynamic definition of callback functions, adding a layer of complexity. This approach makes it difficult to trace the origin and flow of these callbacks, as their definitions and bindings are spread across multiple files and only resolved dynamically.
Challenge#3: Inadequate in Locating WeChat-Specific Components in the UI Widgets Tree. To effectively load subpackages, we need to conduct a detailed UI exploration of the relevant pages. While this process may seem straightforward, the distinctive WXML components introduce unique challenges. These challenges primarily stem from the limitations of traditional testing tools. Designed specifically to align with the WeChat ecosystem, WXML components diverge significantly from the typical Document Object Model (DOM) structures or native UI frameworks familiar to conventional tools. This divergence renders tools like UIAutomator ineffective in identifying and locating these components directly from the GUI widgets tree. This limitation necessitates the development or adaptation of specialized testing tools and methodologies that can cater to the unique structure and behavior of WXML components, ensuring effective testing of MiniApps.
@@ 3.2. Insights for Potential Solutions
Upon evaluating the complexities inherent in MiniApps, particularly considering the aforementioned challenges, we distill three refined insights for potential solutions.
Insight#1: Integrating Cross-Language Data Flow to Constructing Complete UI State Transition Graph. To address Challenge#1, the primary focus is on constructing a detailed and comprehensive topological structure of MiniApps, with an emphasis on enhancing UI state transition analysis through the integration of cross-language data flows. When page routing is implemented through the <navigator> component or platform APIs, the target page paths need to be determined based on data flow analysis to identify the exact values. Especially when dynamic binding syntax is used to specify target paths in WXML, only a thorough data flow analysis of the JavaScript logic can accurately determine the destination page. This multifaceted approach captures the intricate interactions between WXML and JavaScript, ensuring a thorough understanding of data propagation and UI state transitions, thereby providing a robust foundation for page transition analysis and automated UI exploration.
Insight#2: Constructing Complete Cross-Page/Package Callback Control Flow by Analyzing Dynamic Definitions. To tackle Challenge#2, the strategy involves integrating context binding directly into the callback control flow analysis. This approach focuses on analyzing how context (this) propagates and is applied to define the callback functions across different files and packages. By examining how callbacks are defined and where they are bound to specific contexts, we can create a more accurate and comprehensive representation of the callback control flow. Integrating context binding into the control flow provides a comprehensive understanding of the callback functions within a page, enabling better tracking of event-driven behaviors in MiniApps.
Insight#3: Enhancing Existing Dynamic Testing Tools with Fuzzy Matching for Platform-Specific Components. In response to Challenge#3, the strategy involves augmenting existing dynamic testing tools to better handle platform-specific MiniApp widgets. This enhancement would be achieved by implementing a fuzzy matching mechanism based on key component attributes like name, type, and text. By establishing an accurate correlation between static WXML components and dynamic UI Widget Tree elements, we can significantly improve the identification of MiniApp-specific widgets during dynamic runtime. This approach not only ensures the correct recognition of unique MiniApp components but also enhances the overall effectiveness of dynamic testing.
@@ 4. Methodology
In response to the three challenges, we propose MiniScope, an automated UI exploration and privacy inconsistency analysis framework based on two-phase iterative hybrid program analysis, as shown in Figure 3. The framework consists of three key modules: MiniApp Dependency Graph Generator (§4.1), Directed UI Explorer (§4.2), and Privacy Inconsistency Detector (§4.3), which further includes a policy analyzer and a privacy practice monitor, thereby facilitating flow-to-policy cross-validation. Our methodology unfolds in two distinct phases, each involving a round of static analysis followed by a round of UI exploration. The entire workflow is structured as follows:
Main Package Analysis (Phase One). The initial phase of MiniScope is dedicated to dynamically loading subpackages through UI exploration to obtain the complete package. Firstly, MiniScope initializes the MiniApp Dependency Graph (MDG) for the main package. Subsequently, based on the guidance of MDG, the Directed UI Explorer employs a subpackage-directed breadth-first traversal strategy to explore subpackage pages, thus obtaining the complete package of the MiniApp.
Complete Package Analysis (Phase Two). In the second phase, MiniScope tackles the precise identification of privacy practices based on the analysis of the merged complete package. Mirroring the approach of the previous phase, we commence with constructing the complete MDG through static analysis. During the UI exploration stage, MiniScope adopts a privacy-practice-directed depth-first strategy, which is instrumental in conducting a runtime exploration of sensitive behaviors, allowing for a nuanced understanding of the MiniApp’s privacy practices.
@@ 4.1.1. Analyzing UI State Transitions
UTG construction. During UTG construction, MiniScope analyzes the rendering and logic layers separately: (1) For the rendering layer, MiniScope focuses on the <navigator> widgets in the WXML file responsible for page navigation, with the open-type attribute signifying the routing method. (2) For the logic layer, MiniScope chiefly evaluates JavaScript code related to page routing APIs. For instance, wx.navigateTo adds the current page to the page stack and navigates to a new page. By examining five methods in the <navigator> widget and five types of page routing APIs in the logic layer as shown in Table 2, MiniScope can delineate all possible page transitions in the parsing function $\lambda_{U}$ and statically construct the UI state transition graph $G_{U}$. For statically defined page transition paths, MiniScope directly embeds these paths into the UTG. However, for dynamically bound page transitions, where the target page is represented as a variable rather than a hardcoded value, MiniScope records the variable name in the UTG as a placeholder. Subsequently, during data flow analysis, MiniScope resolves the value of such variables by tracking data propagation and value assignments in the code. This two-step approach ensures that dynamically defined transitions are accurately captured and added to the UTG once their target pages are resolved.
Figure 4. A simplified UTG example of Express MiniApp. Corresponding to Figure 2, in Event 1, the user triggers ontap event of <navigator>, leading to a transition from pages/myInfo to pages/takePhoto; in Event 2, the user triggers ontap event of <button text=‘‘manual input’’>, and the corresponding event handler callback function naveToCheckID is executed, naveToCheckID callback function, resulting in a transition from pages/myInfo to pages/checkID.
A UTG example. Figure 4 illustrates a simplified UTG example of the Express MiniApp. In this example, the set $V_{U}$ contains three pages: ‘‘pages/myInfo/index’’(A), ‘‘pages/takePhoto/index’’(B), and ‘‘pages/checkID/index’’(C). In the MiniApp framework, all page widgets must be registered in the WXML file. We determine page states $V_{U}$ by traversing through these widgets. For example, PageA consists of six <navigator> widgets, and when we click on the <navigator url=‘‘/pages/takePhoto/index’’>, it will navigate to the corresponding page, making the widget a transition condition of the set $E_{U}$ from PageA to PageB.
@@ 4.1.2. Analyzing Callback Control Flows
Determining CCFG entry point. Prior to constructing, it is necessary to determine the CCFG entry points which are event-driven callback functions. This is accomplished by performing cross-language analysis on rendering layer widgets and their associated event handler callback functions. User interactions with these widgets trigger the corresponding event handling logic, activating the associated control flows within the MiniApp. As an example shown in Figure 2 (❷), a widget with ‘‘bindtap = takePhoto’’ attribute will bind to the Tap event, and execute the takePhoto function when a Tap event is triggered. Furthermore, the lifecycle callbacks can also serve as CCFG entry points as they are tied to specific lifecycle events, executing relevant logic when these events occur.
CCFG construction. From the identified entry points, MiniScope constructs the rest of the function call chain within the JavaScript language to obtain a complete CCFG. Considering the highly customizable, modular nature of MiniApps, a tailored cross-file/package callback control flow analysis approach is necessary. Function definitions within MiniApp pages resemble class method definitions, as they’re all defined within the Page() object instance. During initialization, developers provide all static-defined lifecycle callback functions, event handler functions, and custom functions to the Page object as attribute methods. Method intercommunication requires the ‘‘this’’ keyword, referring to the current Page object. Notably, some commonly used callback functions can be defined in shared JavaScript files and then dynamically imported across various pages. Leveraging these observations, we design a cross-file callback control flow analysis algorithm based on the JavaScript Abstract Syntax Tree (AST). This algorithm takes into account the unique patterns of function usage and invocation in MiniApps, including the shared use of callback functions across different files and subpackages.
@@
Cross-File Callbacks Resolution Algorithm. The following pseudocode in Algorithm 1 outlines the algorithm designed to identify and track cross-file callback definitions. The algorithm takes inputs including the AST $G_{A}$ of the current page and a set of $bindCalls$ obtained from the WXML. The output of the algorithm is the dynamically defined cross-file callback functions $exCalls$.
Step 1: Identify Imported Modules. The algorithm first identifies all the modules imported into the current file, whether through import statements in ES6 or require calls in CommonJS. This is accomplished by the getimportedModules function. This function iterates over all nodes in the AST $G_{A}$. If a node is an ImportDeclaration, it records the module specified by the import statement (Lines 3-5). If a node is a VariableDeclarator with an initialization that is a CallExpression calling require, it similarly records the module specified (Lines 6-10). The function returns a dictionary importedModules mapping import specifiers to their sources (Line 12).
Step 2: Identify Cross-File Callbacks. With the imported modules identified, the algorithm proceeds to identify cross-file callback functions. This is handled by the getCrossfileCallbacks function. This function iterates over the nodes in $G_{A}$ and looks for CallExpression nodes (Line 16). If a CallExpression contains a ThisExpression in its arguments and the callee is one of the importedModules (Lines 16-17), it retrieves the position of the ThisExpression (Line 18) and the path of the module (Line 19). It then traverses the AST of the imported module to find matching CallExpression nodes and their contexts (Line 21). If it finds a MemberExpression with the correct context (Line 24) and the right child in bindCalls (Line 25), it appends the parent of this expression to exCalls (line 26). Finally, the function returns a node list of cross-file callback functions exCalls.
1
Input: AST $G_{A}$ of the page, a set of $bindCalls$
Output: Dynamically defined cross-file callback functions $exCalls$
2
3
4Function getimportedModules($G_{A}$):
@@
7                   $importedModules\left[node.specifier\right]\leftarrow node.source$
8
@@
10            if $node$ is $VariableDeclarator$ then
11                   if $node.init$ is $CallExpression$ and $node.init.callee.name$ is $require$ then
12                         $importedModules[node.id.name]\leftarrow node.init.arguments[0].value$
13
14                   end if
15
16             end if
17
@@
21
22Function getCrossfileCallbacks($G_{A},importedModules,pageMethods$):
23       for $node\in G_{A}.nodes$ do
@@
25                   if $ThisExpression\in node.arguments$ and $node.callee\in importedModules.keys()$ then
26                         $pos\leftarrow node.arguments.index(ThisExpression)$
27                         $path\leftarrow importedModules\left[node.callee\right]$
28                         for $exnode\in AST(path)$ do
29                               if $exnode$ is $CallExpression$ and $exnode.callee==node.callee$ then
30                                     $ctx\leftarrow exnode.argument.pos$
31
@@
33                              if $exnode$ is $MemberExpression$ and $exnode.leftchild==ctx$ then
34                                     if $exnode.rightchild\in bindCalls$ then
35                                           $exCalls.append(exnode.parent)$
36
@@
40
41                         end for
42
@@
50
Algorithm 1 Cross-File Callbacks Resolution